OpenStudy (anonymous):

solve a-8/a+5=a^2-12/a+5

OpenStudy (lgbasallote):

uhmm where does the problem start?

OpenStudy (anonymous):

a-8/a+5 = a^2-12/a+5

OpenStudy (anonymous):

a=2

OpenStudy (lgbasallote):

\[\large \frac{a-8}{a+5} = \frac{a^2 - 12}{a+5}\] is that right?

OpenStudy (anonymous):

[a = 2], [a = -1/2+(1/2*I)*sqrt(7)], [a = -1/2-(1/2*I)*sqrt(7)]

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

well if you cross multiply... \[\large \frac{\cancel{(x+5)}(a-8)}{\cancel{x+5}} = a^2 - 12\] \[a - 8 = a^2 - 12\] put all of them on one side \[\large a^2 - a - 12 + 8 = 0\] \[\large a^2 -a -4 = 0\]

OpenStudy (lgbasallote):

you can use completing the square or quadratic formula

OpenStudy (lasttccasey):

a = 2 is wrong.. factor, a−8=a2−12 simplify and use quadratic formula to get: \[a = \frac{(1\pm \sqrt{17})}{2}\]